OpenStudy (anonymous):

the ordered pair (-2, ? ) so that it is a solution to the linear equation y = -5x - 4 6 - 14 - 11 3

OpenStudy (anonymous):

use the equation y=-5x-4. Plug x=-2 and solve for y. see what you will get??

OpenStudy (anonymous):

-11?

OpenStudy (anonymous):

y=-5(-2)-4 =10-4 =6 that point is (-2,6)